OpenStudy (anonymous):

Find the distance between the two points and the midpoint of the segment joining them. (Do not use mixed numbers in your answers.) (-4, 3), (-3, 4)

OpenStudy (anonymous):

The Euclidean distance between (x_1,y_1) and (x_2,y_2) is defined as \[ \sqrt{(x_1-x_2)^2+(y_1-y_2)^2} \] The mid point is \[ \left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\right) \] Now just plug in the numbers into the formulas I have given you!